# Bayesian Inference for Consistent Predictions in Overparameterized Nonlinear Regression

### v1

@@ 6.1. Dataset
We apply the proposed method to the ARCENE dataset, which was originally curated for the NIPS 2003 feature selection challenge for the binary classification problem. The dataset consists of mass spectrometric measurements from cancer patients and healthy individuals, with the goal of distinguishing between the two groups based on these measurements. The ARCENE dataset, accessible through the UC Irvine Machine Learning Repository [16], was constructed by merging three separate datasets: National Cancer Institute (NCI) ovarian cancer data; NCI prostate cancer data; and Eastern Virginia Medical School (EVMS) prostate cancer data. In total, the public dataset contains $200$ samples, comprising $88$ cancer samples and $112$ normal samples. These data are categorized as cancer samples and control samples, with $10000$ continuous covariates, with 7000 dimensions corresponding to real mass-spectrometry peaks and $3000$ dimensions representing random noise variables (probes). The primary challenges posed by the dataset lie in its high dimensionality and the presence of overlapping classes. The goal here is to discriminate between cancer and normal samples.
@@ 6.2. Experimental Setting
We compared the performance of five distinct methods.
  • Logistic Regression with the effective spectral (ES) prior distribution: The proposed Bayesian methodology, as described in Section 2.2, was applied to the logistic regression model to make predictions on unseen data.
  • Logistic Regression with the Horseshoe (HS) prior: To handle the high-dimensionality of covariates, we implemented logistic regression with the Horseshoe prior [12]. The implementation was carried out using the ”horseshoenlm” package in R [38].
  • Logistic Regression with the Normal (NM) prior: We investigated the performance of logistic regression with a standard normal prior distribution, which is the most commonly used prior and serves as a ridge-like regularization [15]. This prior allows for flexibility in modeling the coefficient distribution and can be easily implemented using the Pyro probabilistic programming language [3].
  • Logistic Regression with the Laplace (LP) prior: We also explored the performance of logistic regression with a standard Laplace prior distribution, which promotes sparsity in the coefficient estimates as a Bayesian Lasso [43]. This prior has been shown to be effective in high-dimensional settings and can be efficiently implemented using Pyro, leveraging its support for non-Gaussian prior distributions [3].
  • LightGBM: As a baseline method for table data prediction tasks, we implemented LightGBM [27], an efficient and scalable gradient boosting framework. We adopted the ”lightgbm” package in Python for the implementation and trained the model using Bayesian optimization [48, 14] via the ”Optuna” package [1].
For all methods, we conducted performance evaluation using 5-fold cross-validation. In each fold, the model was trained on the training data, and predictions were made on the test data. As in the previous section, the evaluation metrics used to assess predictive performance were Accuracy, Area Under the ROC Curve (AUC), UM, and CC.
@@ 6.3. Results
The results presented in Table 1 provide valuable insights into the performance of the proposed Bayesian logistic regression approach with different prior distributions, as well as its comparison to the LightGBM baseline. Overall, Logistic Regression with the ES prior distribution achieves the highest accuracy and AUC among all the methods, demonstrating its effectiveness in capturing the underlying structure of the high-dimensional ARCENE dataset. The superior performance of the ES prior can be attributed to its ability to adapt to the intrinsic dimensionality of the data by incorporating the empirical spectral information of the covariance matrix [59].
Let us take a closer look at the comparison. The horseshoe prior, known for its flexibility and sparsity-inducing properties [12], also performs well, with the second-highest accuracy and AUC. This suggests that the regression on ARCENE dataset may exhibit some degree of sparsity, which the HS prior is able to capture effectively. However, the ES prior outperforms the HS prior, indicating that the adaptive nature of the ES prior is more beneficial in this context. Also, the fact that the $3000$ covariates are essentially irrelevant to the regression and have zero coefficients but are not highly sparse, may be the reason why the proposed method prevails. The LP and NM priors, while still providing competitive results, do not perform as well as the ES and HS priors. Nevertheless, their performance is still superior to the LightGBM baseline, highlighting the advantages of the Bayesian approach in handling excess covariates.
@@
Table 1. Data Analysis: Data Analysis: Comparison of classification performance between our method, LightGBM. The metrics are Accuracy, Area Under the Curve (AUC), the proportion of uncertain predictions among misclassifications (UM), and the proportion of correct predictions among certain classifications (CC).
@@ 7. Discussion
This study introduces a novel Bayesian approach for prediction in overparameterized generalized linear models, highlighting its theoretical foundations and practical applications. The primary theoretical contribution lies in establishing posterior contraction under appropriate conditions, validating the use of the posterior distribution as a reliable approximation of the true data-generating distribution. Empirical evidence from simulations and real-world datasets corroborates the effectiveness of the proposed method in delivering accurate predictions and reliable uncertainty estimates.
In the realm of Bayesian high-dimensional inference, the proposed approach distinguishes itself by directly addressing non-sparsity while offering theoretical guarantees for prediction with nonlinearity. This unique perspective paves the way for new avenues of Bayesian inference in overparameterized models.
This study marks a significant advance in overparameterization theory by achieving consistency in nonlinear regression problems. The robustness of posterior contraction theory to model complexity (nonlinearity) plays a pivotal role in attaining consistency because it solely necessitates control of the covering number, the decay of the prior distribution in its tail region, and the concentration of the prior distribution around the true value of the parameter.
Finally, we conclude this paper by discussing its limitations and identifying potential areas for improvement. The theoretical analysis hinges on specific assumptions regarding the covariance matrix, which may not always hold in real-world scenarios. Future research could focus on relaxing these assumptions and extending the results to more general settings as deep neural networks generalize to a wide range of data. Moreover, the computational complexity associated with sampling from the posterior distribution can pose challenges for large-scale datasets. Also, further improvements in uncertainty estimation could also be combined with the more flexible approximation such as normalizing flow [44, 45] and Stein Variational Gradient Descent [35].
@@ Acknowledgements
I would like to express my sincere gratitude to my advisor, Masaaki Imaizumi, for his patient and rigorous supervision. This work was supported by JSPS KAKENHI (22J21090) and JST ACT-X (JPMJAX23CS).
@@ Appendix A Proof of the Main Results
In this section, we provide proof of the main results presented in the previous sections. The key is to establish the posterior contraction rate for the squared $L_{2}(P_{\bm{X}})$ norm of the difference between the true and estimated mean functions, i.e., $\mathbb{E}[\{g(\bm{X}^{\top}\widehat{\bm{\beta}})-g(\bm{X}^{\top}\bm{\beta}^{*})\}^{2}]$. Since the link function $g$ is assumed to be Lipschitz continuous, we have the following upper bound (up to a constant factor):
@@
To this end, we invoke a general result on posterior contraction rates from [17]. Let $\mathcal{P}=\mathbb{R}^{p}$ be the parameter space equipped with the distance $d_{\mathcal{P}}$ induced by the $\|\cdot\|_{\Sigma}$ norm, and define a sequence of restricted parameter spaces
@@
Theorem 2.1 of [17] states that the posterior contraction rate in Theorem 1 holds if there exists a sequence $H_{n}\to\infty$ such that the following three conditions are satisfied for any sufficiently large $n$ and some constant $C>0$:
@@
The proof of (26) follows from a similar argument as in [59], which we outline below for completeness.
The extension to the case where $\sigma^{2}$ is unknown is straightforward. If we take $[1/n,n]$ as $\mathcal{P}_{n}$ when we have an inverse Gaussian distribution as a prior distribution for $\sigma^{2}$, the posterior contraction holds because $\mathcal{P}_{n}$ dominates $\mathcal{P}$ as $n$ increases and the remaining part $\mathcal{P}\setminus\mathcal{P}_{n}$ has only the negligible mass of the prior.
Combining the above arguments, we conclude that the conditions of Theorem 2.1 in [17] are satisfied, and hence the posterior contraction in Theorem 1 holds. The explicit characterization of the contraction rate $\varepsilon_{n}$ in terms of the eigenvalues of $\Sigma$ follows from a careful analysis of the interplay between the rate $\rho_{n}$ in Assumption 1, the upper and lower bounds $(L_{\kappa},U_{\kappa})$ in Assumption 2, and the additional assumption on the eigenvalue decay stated in Theorem 1.
@@ Appendix B Consistency as a Distribution and Point Estimator
Following the discussion on posterior distribution contraction, we further explore the convergence of distributions and estimators using various metrics.
@@
Recall that $K(P,P^{\prime})$ is the KL divergence, $V(P,P^{\prime})$ is the KL variation, and $H(P,P^{\prime})$ is the Hellinger distance between models with different parameters.
@@
    For binary regression with the logistic link, we have
@@
As shown in the previous section, the consistency of $\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}$ holds. Consequently, in the above examples, if the distance between distributions can be bounded by constant times the quantity, we obtain the consistency of the posterior distribution to the true data-generating distribution with respect to the KL divergence, KL variation, and Hellinger distance. Furthermore, because the total variation distance and the Hellinger distance induce the same topology, posterior consistency also holds in terms of the total variation distance.
In connection with distributional convergence, we now shift our focus to the convergence of the posterior mean. If the parameter space $\mathcal{P}$ is convex and the distance $d_{\mathcal{P}}$ on $\mathcal{P}$ is bounded and convex, then the posterior mean serves as a consistent estimator ([e.g., Theorem 6.8 in 20]). The assumption of $\mathcal{P}$ being convex is often satisfied, and the boundedness and convexity of the distance hold, for instance, with the total variation distance. Hence, in the above examples, the consistency of the posterior mean is achieved.
@@ Appendix C Computational Method
In this section, we describe the computational method for the proposed approach. We employ stochastic variational inference [19, 21, 4] to approximate the posterior distribution of the model parameters. Since the closed-form expression of the KL divergence is not available, we use the reparameterization trick [28, 46] to enable efficient gradient-based optimization. The reparameterization trick allows us to express the variational distribution in terms of a deterministic function of the variational parameters and a random variable, making the gradients of the variational parameters computable. The optimization part is performed using the Adam algorithm [25].
@@ C.1. Logistic Regression Example
To illustrate the computational procedure, we consider the example of logistic regression. The model components are as follows:
@@
Indeed, the prior distribution is not just a normal distribution but has a support constraint. However, it can be approximated as a normal distribution if $R$ is taken large enough. Also, a procedure to compute the posterior distribution without support constraints and then project it onto the constrained region is justified by [49].
Our goal is to find the optimal variational parameters ${\widetilde{\bm{\mu}}}_{opt}$ and $\widetilde{\Sigma}_{opt}$ that minimize the KL divergence between the variational approximation $q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})$ and the true posterior $p_{pos}$:
@@
To compute the gradients of the KL divergence with respect to the variational parameters, we employ the reparameterization trick. Specifically, we denote by $\widetilde{\bm{\beta}}$ a sample from the variational distribution $q$ and parameterize it as $\widetilde{\bm{\beta}}=\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}$, where $\widetilde{\bm{\epsilon}}$ is a sample from a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}_{p})$ and $\widetilde{\Sigma}_{1/2}$ is a decomposition of $\widetilde{\Sigma}$, i.e., $\widetilde{\Sigma}=\widetilde{\Sigma}_{1/2}\widetilde{\Sigma}_{1/2}^{\top}$. With this reparameterization, the KL divergence can be approximated as an unbiased estimate
@@
We can now compute the gradients of the KL divergence with respect to the variational parameters. The gradient with respect to $\widetilde{\bm{\mu}}$ is given by
@@
The derived gradients can be used in conjunction with the Adam optimizer to update the variational parameters $\widetilde{\bm{\mu}}$ and $\widetilde{\Sigma}_{1/2}$. The optimization procedure iteratively updates the variational parameters until convergence or a maximum number of iterations is reached.

### v2

@@ 6.1. Dataset
We apply the proposed method to the ARCENE dataset, which was originally curated for the NIPS 2003 feature selection challenge for the binary classification problem. The dataset consists of mass spectrometric measurements from cancer patients and healthy individuals, with the goal of distinguishing between the two groups based on these measurements. The ARCENE dataset, accessible through the UC Irvine Machine Learning Repository [17], was constructed by merging three separate datasets: National Cancer Institute (NCI) ovarian cancer data, NCI prostate cancer data, and Eastern Virginia Medical School (EVMS) prostate cancer data. In total, the public dataset contains $200$ samples, comprising $88$ cancer samples and $112$ normal samples. These data are categorized as cancer and control samples, with $10000$ continuous covariates, with 7000 dimensions corresponding to real mass-spectrometry peaks and $3000$ dimensions representing random noise variables (probes). The primary challenges posed by the dataset lie in its high dimensionality and the presence of overlapping classes. The objective here is to discriminate between cancerous and normal samples.
@@ 6.2. Experimental Setting
The performance of the six distinct methods were compared.
  • Logistic Regression with effective spectral (ES) prior distribution: The proposed Bayesian methodology, as described in Section 2.2, was applied to the logistic regression model to make predictions on unseen data.
  • Logistic Regression with the horseshoe (HS) prior: To handle the high-dimensionality of covariates, logistic regression was implemented using the Horseshoe prior [13]. The ”horseshoenlm” package in R [39] was used for the implementation.
  • Logistic Regression with the normal (NM) prior: We investigated the performance of logistic regression with a standard normal prior distribution, which is the most commonly used prior and serves as a ridge-like regularization [16]. This prior provides flexibility in modeling the coefficient distribution and can be easily implemented using the Pyro probabilistic programming language [3].
  • Logistic Regression with the Laplace (LP) prior: the performance of logistic regression was also explored using a standard Laplace prior distribution, which promotes sparsity in the coefficient estimates as a Bayesian Lasso [44]. This prior has proven effective in high-dimensional settings and can be efficiently implemented using Pyro, leveraging its support for non-Gaussian prior distributions [3].
  • LightGBM: As a baseline method for table data prediction tasks, LightGBM [28], which is an efficient and scalable gradient boosting framework, was implemented adopting the ”lightgbm” package in Python for the implementation and training of the model using Bayesian optimization [49, 15] via the ”Optuna” package [1].
  • Principal Component Regression: As a representative method for high-dimensional data, PCR was implemented with the number of bases selected from $1$ to $30$ by 5-fold cross-validation.
For all methods, performance evaluation was conducted using 5-fold cross-validation. In each fold, the model was trained on the training data, and predictions were made on the test data. As in the previous section, the evaluation metrics used to assess predictive performance were Accuracy, AUC, UM, and CC.
@@ 6.3. Results
The results of the experiment are presented in Table 1. The entries represent the average values of the metrics over $5$-fold cross-validation. Overall, among all the methods, logistic regression with the ES prior distribution achieves the highest accuracy and AUC, demonstrating its effectiveness in capturing the underlying structure of the high-dimensional ARCENE dataset. The superior performance of the ES prior can be attributed to its ability to adapt to the intrinsic dimensionality of the data by incorporating the empirical spectral information of the covariance matrix [60].
Let us take a closer look at the comparison. The horseshoe prior, known for its flexibility and sparsity-inducing properties [13], performs well, with the second-highest accuracy and AUC. This suggests that the regression on ARCENE dataset may exhibit some degree of sparsity, which the HS prior is able to capture effectively. However, the ES prior outperforms the HS prior, indicating that the adaptive nature of the ES prior is more beneficial in this context. Also, the fact that the $3000$ covariates are essentially irrelevant to the regression and have zero coefficients but are not highly sparse, may be the reason why the proposed method prevails. The LP and NM priors, while still providing competitive results, do not perform as well as the ES and HS priors. Nevertheless, their performance is still superior to the LightGBM baseline, highlighting the advantages of the Bayesian approach in handling excess covariates. PCR effectively selects relevant components from a large number of covariates although the results are only moderate and do not match those of our method which retains all covariates without reduction.
@@
PCR 0.810 0.878
Table 1. Comparison of classification performance among logistic regressions with effective spectral (ES, our method), horseshoe (HS), Laplace (LP), and normal (NM) priors, LightGBM, and principal component regression (PCR), using accuracy, area under the curve (AUC), proportion of uncertain predictions among misclassifications (UM), and proportion of correct predictions among certain classifications (CC) as metrics.
@@ 7. Discussion
This study introduces a novel Bayesian approach for prediction in overparameterized generalized linear models, providing a theoretical foundation for practical applications. The primary theoretical contribution lies in establishing posterior contraction under appropriate conditions, validating the use of the posterior distribution as a reliable approximation of the true data-generating distribution. Empirical evidence from simulations and real-world datasets corroborates the effectiveness of the proposed method in delivering accurate predictions and reliable uncertainty estimates.
In the field of high-dimensional Bayesian inference, the proposed approach stands out by directly addressing non-sparsity while providing theoretical guarantees for prediction under nonlinearity. This unique perspective paves the way for Bayesian inference in overparameterized models.
This study marks a significant advance in overparameterization theory by achieving consistency in nonlinear regression problems. The robustness of posterior contraction theory to model complexity (nonlinearity) plays a pivotal role in attaining consistency, necessitating control of the covering number, decay of the prior distribution in the tail region, and concentration of the prior distribution around the true value of the parameter.
Finally, the limitations of the paper can help identify potential areas for improvement. The theoretical analysis hinges on specific assumptions regarding the covariance matrix, which may not always hold in real-world scenarios. Future research can focus on relaxing these assumptions to extend the results to more general settings, as deep neural networks generalize to a wide range of data. Moreover, the computational complexity associated with sampling from the posterior distribution can pose challenges for large-scale datasets. Further improvements in uncertainty estimation can also be combined with more flexible approximation approaches, such as normalizing flow [45, 46] and Stein variational gradient descent [36].
Computer Programs
The computer programs used in the paper for the numerical experiments in Section 5 and the application in Section 6 are publicly available on GitHub repository https://github.com/TomWaka/BA-Overparameterized-NonLinReg. The computational algorithm is outlined in the appendix.
@@ Acknowledgements
I would like to express my sincere gratitude to my advisor, Masaaki Imaizumi, for his patient and rigorous supervision. I am also grateful to Shonosuke Sugasawa for his helpful suggestions. This work was supported by JSPS KAKENHI (22J21090) and JST ACT-X (JPMJAX23CS).
@@ Appendix A Proof of the Main Results
In this section, proofs are provided for the main results presented in the previous sections. The key is to establish the posterior contraction rate for the squared $L_{2}(P_{\bm{X}})$ norm of the difference between the true and estimated mean functions, i.e., $\mathbb{E}[\{g(\bm{X}^{\top}\widehat{\bm{\beta}})-g(\bm{X}^{\top}\bm{\beta}^{*})\}^{2}]$. Since the link function $g$ is assumed to be Lipschitz continuous, we have the following upper bound (up to a constant factor):
@@
To this end, we invoke a general result on posterior contraction rates from [18]. Let $\mathcal{P}=\mathbb{R}^{p}$ be the parameter space equipped with distance $d_{\mathcal{P}}$ induced by the $\|\cdot\|_{\Sigma}$ norm, further defining it as a sequence of restricted parameter spaces
@@
Theorem 2.1 of [18] states that the posterior contraction in Theorem 1 holds if there exists a sequence $H_{n}\to\infty$ such that the following three conditions are satisfied for any sufficiently large $n$ and some constant $C>0$:
@@
The proof of (27) follows from a similar argument as in [60].
The extension to the case where $\sigma^{2}$ is unknown, is straightforward. Taking $[1/n,n]$ as $\mathcal{P}_{n}$ in the case of an inverse Gaussian distribution as a prior distribution for $\sigma^{2}$, the posterior contraction holds because $\mathcal{P}_{n}$ dominates $\mathcal{P}$ as $n$ increases and the remaining $\mathcal{P}\setminus\mathcal{P}_{n}$ part has only the negligible mass of the prior.
Combining the above arguments, we conclude that the conditions of Theorem 2.1 in [18] are satisfied, and hence the posterior contraction in Theorem 1 holds. The explicit characterization of the contraction rate $\varepsilon_{n}$ in terms of the eigenvalues of $\Sigma$ follows from a careful analysis of the interplay between the rate $\rho_{n}$ in Assumption 1, upper and lower bounds $(L_{\kappa},U_{\kappa})$ in Assumption 2, and the additional assumption on the eigenvalue decay stated in Theorem 1.
@@ Appendix B Consistency as a Distribution and Point Estimator
Following the discussion on posterior distribution contraction, the convergence of distributions and estimators is further explored using various metrics.
@@
Recall that $K(P,P^{\prime})$ is the KL divergence; $V(P,P^{\prime})$ is the KL variation; and $H(P,P^{\prime})$ is the Hellinger distance between models with different parameters.
@@
    For binary regression with a logistic link, we have
@@
As shown in the previous section, the consistency of $\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}$ holds. Consequently, in the above examples, if the distance between distributions can be bounded by a constant times the quantity, the consistency of the posterior distribution to the true data-generating distribution is obtained with respect to the KL divergence, KL variation, and Hellinger distance. Furthermore, because the total variation and Hellinger distances induce the same topology, posterior consistency also holds in terms of the total variation distance.
In connection with distributional convergence, we now shift our focus to the convergence of the posterior mean. If the parameter space $\mathcal{P}$ is convex and distance $d_{\mathcal{P}}$ on $\mathcal{P}$ is bounded and convex, then the posterior mean serves as a consistent estimator ([e.g., Theorem 6.8 in 21]). The assumption of $\mathcal{P}$ being convex is often satisfied, and the boundedness and convexity of the distance hold for the total variation distance. Hence, in the above examples, the consistency of the posterior mean is achieved.
@@ Appendix C Computational Method
In this section, the computational method for the proposed approach is described. Stochastic variational inference [20, 22, 4] is employed to approximate the posterior distribution of the model parameters. Since the closed-form expression of the KL divergence is not available, a reparameterization trick [29, 47] was used to enable efficient gradient-based optimization. The reparameterization trick allows the variational distribution to be expressed in terms of a deterministic function of the variational parameters and a random variable, making the gradients of the variational parameters computable. The optimization part was performed using the Adam optimizer [26].
@@ C.1. Logistic Regression Example
To illustrate the computational procedure, let us consider an example of logistic regression. The model components are as follows:
@@
Indeed, the prior distribution is not just a normal distribution but has a support constraint. However, it can be approximated as a normal distribution if $R$ is taken large enough. Also, a procedure that computes the posterior distribution without support constraints and then projects it onto the constrained region, was justified by [50].
Our objectove is to find the optimal variational parameters ${\widetilde{\bm{\mu}}}_{opt}$ and $\widetilde{\Sigma}_{opt}$ that would minimize the KL divergence between the variational approximation $q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})$ and the true posterior $p_{pos}$:
@@
To compute the gradients of the KL divergence with respect to the variational parameters, we employ a reparameterization trick. Specifically, denoting by $\widetilde{\bm{\beta}}$ a sample from the variational distribution $q$, we parameterize it as $\widetilde{\bm{\beta}}=\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}$, where $\widetilde{\bm{\epsilon}}$ is a sample from a standard Gaussian distribution $\mathcal{N}(\bm{0};\bm{I}_{p})$ and $\widetilde{\Sigma}_{1/2}$ are obtained from a decomposition of $\widetilde{\Sigma}$, i.e., $\widetilde{\Sigma}=\widetilde{\Sigma}_{1/2}\widetilde{\Sigma}_{1/2}^{\top}$. With this reparameterization, the KL divergence can be approximated as an unbiased estimate
@@
Now, the gradients of the KL divergence can be computed with respect to the variational parameters. The gradient with respect to $\widetilde{\bm{\mu}}$ is given by
@@
The derived gradients can be used in conjunction with the Adam optimizer to update the variational parameters $\widetilde{\bm{\mu}}$ and $\widetilde{\Sigma}_{1/2}$. The optimization procedure iteratively updates the variational parameters until the maximum number of iterations or convergence is reached.